### v2

# InjecGuard: Benchmarking and Mitigating Over-defense in Prompt Injection Guardrail Models

Abstract
Prompt injection attacks pose a critical threat to large language models (LLMs), enabling goal hijacking and data leakage. Prompt guard models, though effective in defense, suffer from over-defense—falsely flagging benign inputs as malicious due to trigger word bias. To address this issue, we introduce NotInject, an evaluation dataset that systematically measures over-defense across various prompt guard models. NotInject contains 339 benign samples enriched with trigger words common in prompt injection attacks, enabling fine-grained evaluation. Our results show that state-of-the-art models suffer from over-defense issues, with accuracy dropping close to random guessing levels (60%). To mitigate this, we propose InjecGuard, a novel prompt guard model that incorporates a new training strategy, Mitigating Over-defense for Free (MOF), which significantly reduces the bias on trigger words. InjecGuard demonstrates state-of-the-art performance on diverse benchmarks including NotInject, surpassing the existing best model by 30.8%, offering a robust and open-source solution for detecting prompt injection attacks. The code and datasets are released at https://github.com/SaFoLab-WISC/InjecGuard.
InjecGuard: Benchmarking and Mitigating Over-defense in Prompt Injection Guardrail Models
Hao Li1,2thanks: These authors contribute equally to this work., Xiaogeng Liu3, 1University of Electronic Science and Technology of China, 2National University of Singapore, 3University of Wisconsin-Madison
1 Introduction
Prompt injection attacks Perez and Ribeiro (2022); Greshake et al. (2023); Liu et al. (2024) represent a serious and emerging threat to the security and integrity of large language models (LLMs) Brown et al. (2020). These attacks exploit the models’ reliance on natural language inputs by inserting malicious or manipulative prompts, leading to undesirable behaviors such as goal hijacking or sensitive data leakage. For instance, a well-known prompt injection technique involves instructing the LLM to “ignore previous instructions”, Branch et al. (2022); Harang (2023a); Perez and Ribeiro (2022); Willison (2022) which can override built-in safeguards and enable the execution of unauthorized actions. As the deployment of LLMs continues to expand across industries, the risk posed by prompt injection attacks becomes more pronounced, demanding effective and efficient defense mechanisms.
Figure 1: Performance comparison of injection detection: We present the average accuracy across benign, malicious, and over-defense cases, plotted against time efficiency. Our method achieves the best performance across performance and efficiency.
To address it, prompt guard models Meta (2024); ProtectAI.com (2024); Deepset (2024b); fmops (2024); LakeraAI (2024a) have recently been proposed as a promising solution. These models work by analyzing the semantic meaning of the input data to detect malicious intent before it reaches the LLM. Unlike LLM models, prompt guard models are lightweight and computationally efficient, as they do not require the high inference cost associated with LLMs. Additionally, these models operate independently, without the need for victim LLM’s responses, further reducing computation costs compared to approaches like LLM guardrails Inan et al. (2023). These attributes make prompt guard models adaptable to various environments, and an attractive solution in scenarios where speed and resource optimization are critical. We
Figure 2: Left: PromptGuard Meta (2024) 1, Right: ProtectAIv2 ProtectAI.com (2024) 2, which is current SotA prompt guard model. We found that existing prompt guard models have an over-defense issue, i.e., misclassify benign inputs as malicious due to an overreliance on specific trigger words, such as “ignore”.
Despite these advantages, we find that existing prompt guard models face a critical limitation: the issue of over-defense. Over-defense arises when models misclassify inputs due to reliance on shortcuts, resulting in false positives where benign inputs are incorrectly flagged as threats. For instance, as shown in Fig. 2, commonly used words such as “ignore” or “cancel” can be in part of harmless sentences but are misclassified as malicious by many existing (even commercial) prompt guard models. Such over-defense problem reduces LLM accessibility, as the prompt guard model may reject legitimate user requests and block access. It can also cause significant disruptions in real-world applications, particularly in interactive systems like virtual assistants Dong et al. (2023) and medical diagnostic tools Thirunavukarasu et al. (2023), where immediate and reliable access is crucial.
To address this issue, we introduce NotInject, an evaluation dataset specifically designed to assess the over-defense issue of existing models. The dataset contains 339 carefully crafted benign inputs, developed using statistical methods on existing benign and attack datasets. The test cases in NotInject contain trigger words commonly found in prompt injection attacks, while still preserving benign intent. We further divide the dataset into three levels of difficulty, based on the number of trigger words present, enabling more fine-grained evaluation. Through systematic evaluations based on NotInject, we demonstrate that current prompt guard models, including current state-of-the-art (SotA) open-source solutions like ProtectAIv2 ProtectAI.com (2024), suffer from significant over-defense issues, with over-defense accuracy falling below 60%, which is close to random guessing (50%).
In addition to the dataset, we also propose a powerful prompt guard model, InjecGuard, which achieves a superior score in both performance and efficiency compared to other guardrail models (see Fig. 1). Since the training approach of existing prompt guard models are all closed-source and training data is not released, our journey starts with curating a comprehensive collection of training datasets with carefully designed data-centric augmentation techniques for addressing the long-tail problem. To further address the over-defense problem, instead of directly finetuning on a specific dataset (e.g, NotInject), which introduces unfair evaluation results, here, we introduce Mitigating Over-defense for Free (MOF), without relying on any specific over-defense datasets. As a result, InjecGuard achieves SotA performance across multiple benchmarks, including NotInject. Evaluation results show that InjecGuard outperforms existing prompt guard models, achieving over 83% average accuracy in detecting benign, malicious, and over-defense inputs, surpassing the open-sourced runner-up prompt guard model by 30.8%. Remarkably, InjecGuard achieves similar performance to GPT-4o OpenAI (2024a), an advanced commercial LLM, while being an lightweight model trained on DeBERTa He et al. (2023). Furthermore, our model reaches this performance using a fully open-source dataset, unlike some existing prompt guard models that rely on closed datasets fmops (2024); Meta (2024); ProtectAI.com (2024); LakeraAI (2024a), further promoting a transparency and open-source academic research environment.
11footnotetext: https://huggingface.co/meta-llama/Prompt-Guard-86M22footnotetext: https://huggingface.co/protectai/deberta-v3-base-prompt-injection-v2
2 Related Works
Prompt injection attacks. Prompt injection attacks have become a critical threat to LLMs. Since LLMs process inputs in natural language, they often face challenges distinguishing between legitimate user commands and external manipulative inputs, leaving them vulnerable. The concept is first identified in research by Perez and Ribeiro (2022), revealing that LLMs could be misled by simple, crafted inputs, resulting in goal hijacking and prompt leakage. Several studies have been proposed Greshake et al. (2023); Wang et al. (2023); Pedro et al. (2023); Yan et al. (2023); Yu et al. (2023); Salem et al. (2023); Yip et al. (2024); Zhan et al. (2024); Liu et al. (2024); Pasquini et al. (2024); Shi et al. (2024), addressing various aspects of prompt injection attacks, such as handcrafted methods Toyer et al. (2023), automatic attack algorithms Liu et al. (2024), and benchmarks Liu et al. (2023); Debenedetti et al. (2024). Public discussions Harang (2023b); Willison (2022, 2023) have also underscored the risks of prompt injection attacks on commercial LLMs. Prompt Injection datasets are also introduced, such as PINT LakeraAI (2024b), Safeguard-Injection Erdogan et al. (2024), TaskTracker Abdelnabi et al. (2024), and BIPIA Yi et al. (2023), etc.
Prompt guard models. Prompt guard models offer a promising approach to defend against prompt injection attacks by detecting malicious intent in inputs. These methods are computationally efficient and, unlike LLM guardrails Inan et al. (2023), do not require an additional round of victim LLM inference to generate a response. Several prompt guard models have been proposed recently. Fmops fmops (2024) is a guardrail trained by FmopsAI, utilizing DistilBERT Sanh et al. (2020) as the backbone. Deepset Deepset (2024b), PromptGuard Meta (2024), and ProtectAIv2 ProtectAI.com (2024) all employ DeBERTaV3-base He et al. (2023) as the backbone. Deepset Deepset (2024b) is trained on Deepset-Injection Deepset (2024a), a small open-source dataset with 546 samples. PromptGuard is a model trained by Meta with a closed-source dataset. ProtectAIv2 is also trained with a closed-source dataset by Protect AI, achieving SotA performance among open-source guard models Deepset (2024b); fmops (2024); Meta (2024); ProtectAI.com (2024) on PINT LakeraAI (2024b). Furthermore, LakeraGuard LakeraAI (2024a) is a commercial guard model developed by LakeraAI, with undisclosed model details, such as scale and architecture, and available solely through an API.
However, most of these prompt guard models fmops (2024); Deepset (2024b); Meta (2024); ProtectAI.com (2024) suffer from over-defense issues, where they rely on shortcuts triggered by certain keywords to make predictions, leading them to incorrectly categorize a benign input that contains the keywords as malicious. Moreover, most of these models fmops (2024); Meta (2024); ProtectAI.com (2024); LakeraAI (2024a) are trained using closed-source data with undisclosed implementation or training details. In this paper, we introduce an over-defense dataset to systematically evaluate the over-defense issue, and propose a novel and completely open-source training approach that achieves SotA performance across diverse benchmarks, including the over-defense dataset.
Figure 3: Visualization of attention weight. Given an instruction of “[CLS] Can I ignore this warning appeared in my code? [SEP]”, ProtectAIv2 ProtectAI.com (2024) assigns extremely high attention weights to the word “ignore,” leading to misclassification as Injection. In contrast, our method distributes attention across the entire sentence, successfully predicting it as benign.
3 Over-defense Dataset: NotInject
Figure 4: The pipeline for constructing NotInject dataset.
3.1 The Over-defense Issue
While prompt guards offer several advantages, such as low overhead as previously mentioned, we have identified a critical limitation: they exhibit severe over-defense issues, even in some advanced models Meta (2024); ProtectAI.com (2024). Specifically, these models learn a shortcut from certain trigger words, like “ignore”, directly to the final prediction. As shown in Fig. 3 with red lines, ProtectAIv2 ProtectAI.com (2024), one of the SotA prompt guard models on PINT benchmark, distributes excessive and unbalanced attention to the word “ignore”, leading it to incorrectly categorize a benign input as malicious. To further illustrate, we input a benign sentence containing the word “ignore” into two advanced prompt guard models: PromptGuard Meta (2024) and ProtectAIv2 ProtectAI.com (2024), and present the results in Fig. 2. Both models incorrectly classify the benign instruction containing the word “ignore” as malicious.
In this paper, we define the over-defense issue in prompt guard models as the tendency to predict malicious labels when benign sentences contain certain trigger words commonly used in prompt injection attacks.
3.2 Construction of NotInject
To enhance the community’s ability to address the above issues, we introduce an over-defense evaluation dataset that supports systematically evaluating the over-defense issue inherent in prompt guard models. As shown in Fig. 4, to build NotInject, there are three main steps: 1) Trigger word identification, which aims to find candidate trigger words likely to cause over-defense; 2) Trigger word refinement, which filters out misidentified words from the first step; and 3) Corpus generation, which uses LLMs to generate over-defense test cases based on the identified trigger words.
Trigger Words Identification. We begin by collecting two primary datasets: a dataset containing known prompt injection attack examples, denoted as $D_{m}$ (malicious dataset), and a benign dataset comprising typical user inputs devoid of malicious intent, denoted as $D_{b}$. We collect these data from several open-source datasets, such as Alpaca Taori et al. (2023) and Safeguard-Injection Erdogan et al. (2024) (detailed information is provided in the Appendix. A). The injection dataset includes various malicious inputs crafted to manipulate LLMs, whereas the benign dataset represents non-harmful user interactions. Then, as illustrated in Fig. 4, we perform a word frequency analysis on both datasets. For each dataset, we compute the word frequencies to obtain frequency lists $F_{b}$ and $F_{m}$ for the benign and malicious datasets, respectively. We rank these words based on their frequencies from highest to lowest, resulting in two separate lists sorted by occurrence rates. Next, to identify injection-specific words, we compare the word frequency rank between the two datasets. By calculating the rank difference through, $\Delta r(w)=R_{b}(w)-R_{m}(w)$, we recognize words that are more frequent in the injection dataset but less common in the benign dataset. Words that exhibit a significantly higher frequency in the injection dataset are flagged as potential trigger words associated with prompt injections. Our detailed algorithm is shown in Alg. 1. We also visualize the top 20 words identified as trigger words by our method in Fig. 5.
Figure 5: The top-20 rank differences between benign and malicious datasets.
Algorithm 1 Trigger Words Identification
1:Benign dataset $D_{b}$, Malicious dataset $D_{m}$, Integer $k$
2:Compute word frequencies in $D_{b}$ to get frequency list $F_{b}$
3:Compute word frequencies in $D_{m}$ to get frequency list $F_{m}$
4:Sort $F_{b}$ in descending order to get rank list $R_{b}$
5:Sort $F_{m}$ in descending order to get rank list $R_{m}$
6:for all words $w$ in $R_{b}\cup R_{m}$ do
7:     Compute rank difference $(\Delta r(w)=R_{b}(w)-R_{m}(w))$
8:     Add $(w,\Delta r(w))$ to list $L$
9:end for
10:Sort list $L$ in descending order based on $\Delta f(w)$
11:return Top $k$ words from list $L$
Trigger Words Refinement. Recognizing that automated methods may include irrelevant or common words not indicative of prompt injection attempts, we proceed with a word rectification process. We first employ LLM to automatically filter the list of candidate words to remove any that may not be pertinent. Specifically, We assess the potential harmfulness of words by asking GPT-4o-mini (OpenAI (2024b), we use the 2024-07-18 version by default unless otherwise specified) questions like, “Do you think the word of {word} is especially frequent in malicious or prompt attack scenarios?” 1 Next, we perform a manual verification step, where we review the remaining words after LLM filtering to ensure that any words unrelated to prompt injection attacks are removed. Through these two steps, we have compiled a word list with the highest potential for malicious use.
Corpus Generation. With a refined list of trigger words, we then generate sample sentences. We instruct the GPT-4o-mini to craft new sentences that must include a specified number of the selected trigger words 2. Specifically, we select the generated 113 trigger words and create three subsets with distinct difficulty levels, determined by the number of trigger words used in the generation of new sentences. Namely, three subsets separately containing 1, 2, and 3 trigger words are built with 113 benign sentences per subset. It is imperative that these sentences are contextually coherent, semantically meaningful, and do not contain any prompt injection instructions or malicious content. The goal is to represent benign usage of the trigger words in everyday language, ensuring that the sentences reflect natural and diverse linguistic patterns. To accomplish this, we design a carefully curated prompt (see Appendix. C.1) to guide the LLM in generating safe and natural sentences. We then implement a polish process to further ensure the safety of the generated samples. In this process, the generated sentences combined with the prompt in Appendix. C.2 are re-input into the LLM to identify any potential injection vulnerabilities. Following this, we conduct a manual review to confirm the safety of all sentences and report the error ratio of three subsets in Fig. 8. This multi-step refinement process guarantees that all generated sentences are harmless. The final output forms the proposed NotInject, which contains a total of 339 generated samples with 113 for one-word subset, 113 for two-word subset, and 113 for three-word subset. NotInject encompasses a diverse set of topics to enable a thorough evaluation, including common queries from daily life, technique queries (eg., programming, system), virtual creations, and multilingual queries (eg., Chinese, Russian, etc.). The detailed category distribution is presented in Table. 1.
Category one-word two-word three-word
Common Queries 58 49 19
Techniques 16 30 41
Virtual Creation 14 4 24
Multilingual 25 30 29
Table 1: Topic category distribution of NotInject.
footnotetext: 1, 2 Detailed prompts are in the Appendix. C
3.3 Evaluations on NotInject
Here, we present our evaluations based on the proposed NotInject, assessing $5$ existing prompt guard models and an advanced LLM-based guardrail. To provide a comprehensive evaluation, in addition to the over-defense evaluation, we employ a three-dimensional metric. First, we measure malicious accuracy, which reflects how effectively the models detect malicious inputs using attack data from the PINT LakeraAI (2024b) and BIPIA datasets Yi et al. (2023). Second, we evaluate benign accuracy, i.e., how accurately the models classify benign data as non-malicious, using benign data from both the PINT and WildGuard benchmark Han et al. (2024). Finally, we assess over-defense accuracy, which captures the models’ performance when encountering benign inputs that contain trigger words like “ignore”, using our proposed NotInject dataset.
Figure 6: Comparison on 3 dimensions of benign, malicious, and over-defense accuracy.
Based on above, we evaluate existing prompt guard models (i.e., Deepset Deepset (2024b), Fmops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024), LakeraGuard LakeraAI (2024a)) and an advanced LLM-based guardrail designed to detect malicious content (i.e., LlamaGuard3 Dubey et al. (2024)).
The results are shown in Fig. 6, where the x, y, and z axes represent malicious accuracy, benign accuracy, and over-defense accuracy, respectively. As illustrated, none of the existing prompt guard models (i.e., Deepset Deepset (2024b), Fmops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024)), LakeraGuard LakeraAI (2024a)) nor even the more powerful LLM-based guardrail model (i.e., LlamaGuard3 Dubey et al. (2024)) can achieve high accuracy across all three dimensions simultaneously. More importantly, none of the existing open-source prompt guard models Deepset (2024b); fmops (2024); Meta (2024); ProtectAI.com (2024) achieve an over-defense accuracy greater than $60\%$, where $50\%$ represents random guessing. These findings underscore that our NotInject poses a significant challenge for current prompt guard models, effectively revealing the prevalent over-defense issue. Although LakeraGuard LakeraAI (2024a) and LlamaGuard3 Dubey et al. (2024) achieve remarkable performance in over-defense accuracy, their reliability is still limited by suboptimal malicious accuracy. This highlights the urgent need to develop a robust prompt guard model that excels in malicious accuracy, benign accuracy, and over-defense accuracy, with particular emphasis on addressing the over-defense issue.
4 InjecGuard
In this section, we aim to develop a robust prompt guard model that excels in malicious accuracy, benign accuracy, and over-defense accuracy. We also want to emphasize that existing prompt guard models are typically trained in a closed-source manner and disclose their training data. Our goal is to train our model using only publicly available data, thereby promoting transparency and facilitating academic research. Here, we introduce InjecGuard, a prompt guard model that achieves state-of-the-art performance across various benchmarks. To build InjecGuard, we first curate a comprehensive collection of open-source datasets for training and apply data-centric augmentation techniques to address the long-tail problem within the dataset. More importantly, we propose a novel training strategy, MOF, which effectively mitigates over-defense issues without relying on specific over-defense datasets.
4.1 Data Collection and Augmentation
Similar to existing models, we first collect a wide range of prompt injection and benign corpus as the training data. Specifically, we select $20$ open-source datasets, including the benign datasets like Alpaca Taori et al. (2023), the prompt injection datasets like Safeguard-Injection Erdogan et al. (2024), and the datasets generated by existing injection attack, such as TaskTrack Abdelnabi et al. (2024)). However, upon analyzing these datasets, we identify a long-tail issue in these datasets: certain input formats frequently exploited in prompt injection attacks—such as CSV files—are underrepresented. To address this, we implement a data-centric augmentation procedure, generating additional data for these long-tail formats. Using GPT-4o-mini, we create some prompt injection samples in 17 formats, including Email, Document, Chat Conversation, JSON, Code, Markdown, HTML, URL, Base64, Table, XML, CSV, Config File, Log File, Image Link, Translation, and Website. After augmentation, our final training dataset for InjecGuard comprises $61,089$ benign samples and $15,666$ prompt injection samples, where $435$ samples are generated by our data-centric augmentation procedure. Detailed statistics of the training data and prompt for data-centric augmentation are provided in the Appendix. A and C.3, respectively.
4.2 Mitigating Over-defense for Free (MOF)
To address the over-defense issue, we propose Mitigating Over-defense for Free (MOF), a novel method to identify and mitigate such biases, without relying on any specific over-defense datasets.
After training our model with the aforementioned data (Sec. 4.1) using standard supervised learning, we perform a “token-wised recheck” to identify biases in the trained model. This procedure takes every token in the tokenizer vocabulary and inputs each token individually into the trained model. Ideally, these tokens should all be considered “benign” since they are individual tokens without any intent of prompt injection attack. By doing this, we can identify any biased words that are incorrectly predicted as “attack” by the model. We consider the bias toward these tokens as the root cause of the over-defense issue in the model.
After identifying the biased tokens, we prompt GPT-4o-mini to generate benign data (see Appendix. C.1) using random combinations of these tokens (including one-token, two-token, and three-token settings, as mentioned in Sec. 3.2). We generate 1,000 benign samples using this method. Afterward, a LLM refinement process similar to the method mentioned in Sec. 3.2 is employed to ensure the harmlessness of generated data. Subsequently, we incorporate these generated data into the training data introduced in Sec. 4.1 to form the final training dataset. Using this dataset, we retrain our prompt guard model from scratch and get the final InjecGuard.
As illustrated in Fig. 3, the retrained model focuses and makes the final prediction on the overall meaning of the entire input, rather than certain trigger words. And results in Fig. 6 show that our model achieves high accuracy across all three dimensions simultaneously. We also present ablation studies of the aforementioned training design in Tab. 3. More detailed evaluation results are provided in the following section.
5 Evaluations
Performance Overhead
Category Model Over-defense (%) Benign (%) Malicious (%) Average (%) GFLOPs Inference (ms) Efficiency
Prompt Guard Model Fmops fmops (2024) 5.60 34.63 93.50 44.58 24.19 4.43 10.06
Deepset Deepset (2024b) 5.31 34.06 91.50 43.62 60.45 15.22 2.87
PromptGuard Meta (2024) 0.88 26.82 97.10 41.60 60.45 15.28 2.72
ProtectAIv2 ProtectAI.com (2024) 56.64 86.20 48.60 63.81 60.45 15.77 4.05
LakeraGuard LakeraAI (2024a) 87.61 90.89 53.19 77.23 - 710.41 0.11
Large Language Model GPT-4o OpenAI (2024a) 86.73 90.78 79.10 85.53 - 7907.18 0.01
Llama-2-chat Touvron et al. (2023) 76.40 61.03 31.09 56.17 1387.49 3111.36 0.02
LlamaGuard3 Dubey et al. (2024) 99.71 95.18 28.28 74.39 1418.38 787.48 0.09
Prompt Guard Model InjecGuard (Ours) 87.32 85.74 77.39 83.48 60.45 15.34 5.44
Table 2: Performance and overhead comparison between our model, InjecGuard, and other baseline models. InjecGuard surpasses the runner-up (ProtectAIv2) by 30.8% in terms of average accuracy. Additionally, InjecGuard achieves performance comparable to GPT-4o, a commercial LLM, despite being an lightweight model trained on open-source data. We calculate the Efficiency = Average Accuracy/Inference Time.
5.1 Experimental Setups
Evaluation Datasets. As mentioned in Sec. 3.3, we conduct our evaluations based on three aspects: benign accuracy, malicious accuracy, and over-defense accuracy. To fully verify the practicality of our proposed InjecGuard, we ensure that all evaluation datasets are not included in the training data. Specifically, For evaluating the benign accuracy, we utilize benign data from both the PINT benchmark LakeraAI (2024b) and WildGuard benchmark Han et al. (2024). For evaluating the malicious accuracy, We use attack data from both the PINT benchmark LakeraAI (2024b) and BIPIA datasets Yi et al. (2023). For evaluating the over-defense accuracy, we use the proposed NotInject dataset.
Metrics. Since prompt guard models function as text classification systems that predict whether an input text is benign or malicious, we evaluate their performance using Accuracy, calculated as the proportion of correct predictions over the total number of test cases in the evaluation dataset:
Accuracy=Number of Correct PredictionsTotal Number of Test $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Test Cases}}$ (1)
Additionally, we report the computational overhead in terms of Giga Floating Point Operations (GFLOPs), which quantifies the total number of floating-point operations required during inference. GFLOPs provide an estimate of the computational resources needed by the model. We also measure the inference time to assess the computational overhead introduced by the models.
Training Details of InjecGuard. We employ DeBERTaV3-base (He et al., 2023) as the backbone of InjecGuard and train it with a batch size of 32 for 3 epochs, using Adam Diederik (2015) optimizer and linear scheduler. The initial learning rate is set to 2e-5, with a 100-step warm-up phase. Besides, the maximum token length is set to 512.
Baselines. We employ five existing prompt guard models as baselines: Fmops fmops (2024), Deepset Deepset (2024b), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024), and LakeraGuard LakeraAI (2024a), which have been introduced in Sec. 2. We also consider LLM-based methods, including LlamaGuard3 Dubey et al. (2024), Llama-2-chat-7b Touvron et al. (2023), and GPT-4o OpenAI (2024a), which are evaluated by prompting them to determine a given input constitutes a prompt injection attack.
5.2 Main Results
We compare our method with other baselines in terms of performance and overhead. The results are shown in Tab. 2 (detailed results for each subsets deferred to Appendix. B.1).
Performance Comparison. Our InjecGuard demonstrates superior performance compared to existing prompt guard models and even rivals commercial LLMs like GPT-4o. Specifically, InjecGuard achieves an average accuracy of 83.48%, the best-performing commercial prompt guard model, by 6.25%, and exceeding the top open-source model, ProtectAIv2, by 30.83%. Our model excels across all evaluation categories, attaining an over-defense accuracy of 87.32%, benign accuracy of 85.74%, and malicious accuracy of 77.39%. This balanced performance indicates that InjecGuard is highly effective at correctly identifying benign and malicious inputs while minimizing false positives and negatives. The fact that InjecGuard achieves results comparable to GPT-4o, despite being based on the open-source data and lightweight backbone, highlights the efficiency and effectiveness of our training approach. Notably, according to Tab. 2, the Efficiency (Performance/Inference time) of our model is 503 times higher than that of GPT-4o.
The results also prove that existing prompt guard models suffer from overwhelming issues with over-defense, as reflected by their low over-defense accuracy scores. Models like Fmops, Deepset, and PromptGuard exhibit over-defense accuracies as low as 5.60%, 5.31%, and 0.88%, respectively, indicating a tendency to incorrectly classify benign inputs that have trigger words as malicious. And ProtectAIv2 has an over-defense accuracy of 56.64%, which is close to random guessing (50.00%). This issue underscores the challenges posed by our proposed over-defense dataset. In addition, the superior over-defense accuracy of InjecGuard surpasses the previously best open-source prompt guard model (ProtectAIv2) by 54.17%, showing that our model effectively handles the challenging cases presented by the dataset, and confirming the efficacy of our novel training approach MOF in mitigating over-defense issues prevalent in existing models. Note that MOF does not require any over-defense dataset for training; instead, it automatically generates adaptive training data, verifying it as both practical and highly effective.
Overhead Comparison. Our model, InjecGuard, excels in both computational efficiency and performance compared to existing baseline models. As shown in Table 2, InjecGuard  achieves an average accuracy of 83.48% while maintaining a GFLOPS count of 60.45 and an inference time of only 15.34 milliseconds. With an efficiency score of 5.44 (performance divided by inference time), InjecGuard delivers robust results swiftly and resource-effectively, making it ideal for environments prioritizing both accuracy and efficiency. In contrast, LLMs like GPT-4o achieve slightly higher accuracy (85.53%) but at a significant computational cost, with an inference time of 7907.18 milliseconds and an efficiency score of 0.01. Similarly, models such as Llama-2-chat and LlamaGuard3 require thousands of GFLOPS and longer inference times, resulting in much lower efficiency. This disparity highlights the impracticality of deploying these LLMs in scenarios where computational resources and quick response times are essential. Therefore, InjecGuard presents a highly effective and efficient alternative for real-world applications demanding both high accuracy and operational efficiency.
Training Design Over-defense (%) Benign (%) Malicious (%) Average (%)
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + MOF with Retraining from Scratch (Sec. 4.2) 89.38 84.73 71.57 81.89
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
   + MOF with Finetuning (Sec. 4.2) 68.14 82.11 74.81 75.02
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
   + MOF with Retraining from Scratch (Sec. 4.2) 87.32 85.74 77.39 83.48
Table 3: Ablation study of each training design introduced in Sec. 4.
5.3 Ablation Studies
The ablation study presented in Tab. 3 offers key insights into the effects of our training components on InjecGuard performance. Starting with the Basic Dataset, the model establishes a baseline average accuracy of 74.64%, with over-defense, benign, and malicious accuracies of 75.22%, 78.53%, and 70.17%, respectively. Introducing Data-centric Augmentation alone leads to a noticeable improvement in benign accuracy (from 78.53% to 81.36%) and malicious accuracy (from 70.17% to 75.95%), enhancing the model’s ability to correctly classify both benign and malicious inputs. However, this augmentation comes at a cost, as evidenced by a significant reduction in over-defense accuracy from 75.22% to 64.31%. This decline indicates that while data-centric augmentation enriches the training data and improves classification capabilities, it inadvertently exacerbates the over-defense issue, making the model more prone to incorrectly classifying benign inputs as malicious—a challenge commonly observed in existing models.
The introduction of the MOF, particularly when combined with Retraining from Scratch, addresses the adverse effects of data-centric augmentation on over-defense accuracy. Applying MOF with Retraining from Scratch to the Basic Dataset lifts the average accuracy to 81.89%, with over-defense, benign, and malicious accuracies at 89.38%, 84.73%, and 71.57%, respectively. This substantial improvement highlights MOF’s effectiveness in mitigating over-defense issues caused by data-centric augmentation. As previously mentioned, employing Data-centric Augmentation can potentially reduce the performance of overdefense. However, when combined with MOF and retraining from scratch, the model achieves its highest average accuracy of 83.48%, with over-defense, benign, and malicious accuracies of 87.32%, 85.74%, and 77.39%. This combination leverages the strengths of both data augmentation and mitigation strategies, resulting in a robust model that not only benefits from enhanced classification performance but also maintains high over-defense accuracy. The ablation study demonstrates that data-centric augmentation alone can improve certain aspects of model performance, and the integration of MOF techniques is essential to counterbalance the increased over-defense, thereby ensuring that InjecGuard achieves optimal and balanced performance across all metrics.
6 Conclusions
In this study, we delved into the over-defense phenomenon observed in existing injection-based prompt guard models. To address this issue, we proposed a novel over-defense data sampling framework and introduced the NotInject benchmark to systematically evaluate the extent of over-defense in target prompt guard models. Furthermore, we presented InjecAgent, an advanced prompt guard model trained using a novel robust training strategy, MOF, designed to effectively mitigate model bias toward specific words. Our experimental results on various benchmark demonstrate that InjecAgent significantly outperforms existing prompt guard models across multiple dimensions, with substantial improvements in performance and robustness. To the best of our knowledge, this is the first work to provide a fully open-source prompt guard model against injection, including the training dataset, strategies, code, and model. We believe this will further promote transparency and foster an open-source academic research environment for advancing future LLM safety exploration.
Limitations
While our work shows significant improvement in mitigating over-defense in prompt guard models, the NotInject dataset, while carefully designed, may not fully capture the diversity of real-world benign inputs, particularly in domain-specific applications. This could result in the underestimation of models’ over-defense tendency in complex, sensitive fields such as healthcare or finance. However, as our comprehensive evaluations have shown, the current design of NotInject is sufficient to reveal the over-defense issue in existing prompt guard models, highlighting the urgent need for improved approaches in this community. To further enhance the diversity of NotInject, our future work will incorporate domain-specific data through collaboration with industry partners.
Ethics Statement
We are committed to advancing the security and integrity of LLMs responsibly. In this research, we introduce NotInject, a dataset designed to assess and mitigate the over-defense issue in prompt guard models. Additionally, we introduce InjecGuard, a powerful prompt guard model developed using our novel approach, MOF, aimed at enhancing LLM security. All data used are synthetically generated or sourced from publicly available datasets, ensuring that no personal or sensitive information is involved. This approach safeguards privacy and complies with ethical standards regarding data use.
While our work focuses on enhancing defensive mechanisms against prompt injection attacks, we acknowledge the potential for dual use in security research. We encourage the ethical and responsible use of NotInject to improve LLM security and not for malicious purposes. By addressing over-defense, we aim to reduce false positives and enhance accessibility for all users when prompt guard models are deployed. Our commitment to transparency is reflected in making both the dataset and model fully open-source, fostering collaboration, and allowing others to verify, replicate, and build upon our work for the betterment of the field.
Appendix
Appendix A Datasets
A.1 Benign dataset
Our benign data are collected from 14 open-source datasets, and our augmented over-defense dataset in Sec. 4.1. The open-source datasets include Alpaca Taori et al. (2023), chatbot_instruction_prompts Palla (2024), open-instruct VMware (2023), xstest-v2-copy Röttger et al. (2023), grok-conversation-harmless HuggingfaceH4 (2023), prompt-injections Deepset (2024a), safe-guard-prompt-injection Erdogan et al. (2024), awesome-chatgpt-prompts Akın (2023), no_robots Rajani et al. (2023), ultrachat_200k Ding et al. (2023), TaskTracker Abdelnabi et al. (2024), BIPIA_train Yi et al. (2023), jailbreak-classification Hao (2023), and Question Set Shen et al. (2023). The data distribution is shown in Tab. 4.
A.2 Malicious dataset
Our malicious data are built based on 12 open-source datasets, and our augmented dataset in Sec. 4.1. The open-source datasets include InjecAgent Zhan et al. (2024), prompt-injections Deepset (2024a), hackaprompt-dataset Schulhoff et al. (2023), safe-guard-prompt-injection Erdogan et al. (2024), ChatGPT-Jailbreak-Prompts Romero (2023), vigil-jailbreak-ada-002 Swanda (2023), Prompt-Injection-Mixed-Techniques Yugen.ai (2023), TaskTracker Abdelnabi et al. (2024), StruQ Chen et al. (2024), BIPIA_train Yi et al. (2023), jailbreak-classification Hao (2023), and Question Set Shen et al. (2023). The data distribution is shown in Tab. 5. Our augmented dataset consists of a large number of long-tail data types, such as XML, HTML, Markdown, etc. The specific data type distribution is shown in Tab. 6.
Source Count
Alpaca 4,000
chatbot_instruction_prompts 16,000
open-instruct 12,000
xstest-v2-copy 450
grok-conversation-harmless 4,000
prompt-injections 343
safe-guard-prompt-injection 5,740
awesome-chatgpt-prompts 170
no_robots 1,500
ultrachat_200k 3,000
TaskTracker 11,386
BIPIA_train 558
jailbreak-classification 517
Question Set 643
over-defense augmented set 762
Table 4: Benign Dataset Source.
Source Count
InjecAgent 111
prompt-injections 203
hackaprompt-dataset 5,000
safe-guard-prompt-injection 2,496
ChatGPT-Jailbreak-Prompts 79
vigil-jailbreak-ada-002 104
Prompt-Injection-Mixed-Techniques 1,174
TaskTracker 3,316
StruQ 20
BIPIA_train 558
jailbreak-classification 527
Question Set 1,643
LLM Augmented set 435
Table 5: Injection Dataset Source.
Category Count
Email Injection 48
Document Injection 25
Chat Conversation Injection 25
JSON Injection 23
Code Injection 23
Markdown Injection 23
HTML Injection 23
URL Injection 23
Base64 Injection 23
Table Injection 23
XML Injection 23
CSV Injection 23
Config File Injection 23
Log File Injection 23
Image Link Injection 23
Translation Injection 27
Website Injection 34
Table 6: Categories of LLM augmentated set.
Method NotInject Wildguard Pint BIPIA
one-word two-word three-word Benign Benign Injection overall Injection
Deepset 5.31 7.08 3.54 50.98 17.13 98.32 57.73 84.67
Fmops 5.31 5.31 6.19 50.88 18.38 98.32 58.35 88.67
PromptGuard 2.65 0.00 0.00 6.69 46.94 94.19 70.56 100.00
ProtectAIv2 76.11 47.79 46.02 75.18 97.22 88.53 92.88 8.67
LakeraGuard 90.27 92.92 79.64 82.60 99.18 96.38 97.78 12.00
GPT-4o 95.58 85.84 78.76 84.24 97.31 92.19 94.75 66.00
Llama-2-chat 82.30 79.65 67.26 74.07 47.99 21.84 34.92 40.34
LlamaGuard3 100.00 100.00 99.12 95.16 95.19 16.89 56.04 39.67
InjecGuard (Ours) 91.15 89.38 81.42 76.11 95.36 86.43 90.90 68.34
Table 7: Full results of comparison between existing injection guardrails.
Appendix B Additional Experimental Results
B.1 Full Results
In Tab. 2, we have illustrated the comprehensive results on different dimensions, such as over-defense, benign, and malicious. In this section, we present detailed results for each evaluation benchmark across all dimensions, including our NotInject, Wildguard Han et al. (2024), PINT-benchmark LakeraAI (2024b), and BIPIA Yi et al. (2023). The results are shown in Tab. 7.
B.2 Visualization of NotInject dataset
To further investigate the advantages of our method on confronting over-defense, we select the benign sentence from our over-defense dataset and perform a visualization to conduct qualitative analysis for model predictions. The results presented in Fig. 7 reveal that although the input is entirely safe, both PromptGuard Meta (2024) and ProtectAIv2 ProtectAI.com (2024) predict it as an injection with high confidence.
In contrast, our InjecGuardaccurately classifies the input as safe, highlighting the efficacy of the over-defense dataset in evaluating over-defense tendencies and demonstrating the robustness of the proposed InjecGuard.
Figure 7: Different guard models’ prediction visualization.
Appendix C Prompts
In this section, we illustrate the prompts used in our method.
C.1 Word-based Generation Prompt
In Sec. 3.2 and Sec. 4.2, we leverage LLMs to generate benign sentences based on trigger words or tokens for both the NotInject dataset and our MOF strategy. The prompts used in this process are detailed in Fig. 9.
C.2 Refinement Prompt
In Sec. 3.2 and Sec. 4.2, we utilize LLMs to conduct malicious content detection on trigger-based generated sentences, ensuring the harmlessness of the generated content. In our experiments, the same prompts are employed to facilitate malicious content detection by GPT-4o OpenAI (2024a) and Llama-2-chat Touvron et al. (2023). The detailed prompts are provided in Fig. 10.
C.3 Long-tail Format Augmentation Prompt
In Sec. 4.1, to address the issue of long-tail distribution in certain data formats (such as CSV, XML) within the current dataset, we propose a data augmentation approach leveraging LLM. The detailed augmentation prompt is illustrated in Fig. 11.
Figure 8: The error ratio of three subsets after LLM generation.
Figure 9: Word-based Generation Prompt.
Figure 10: LLM Injection Detection Prompt.
Figure 11: LLM Augmentation Prompt.

### v3

@@
 InjecGuard demonstrates state-of-the-art performance on diverse benchmarks including NotInject, surpassing the existing best model by 30.8%, offering a robust and open-source solution for detecting prompt injection attacks.
-The code and datasets are released at https://github.com/SaFoLab-WISC/InjecGuard.
+The code and datasets are released at https://github.com/leolee99/InjecGuard.
 InjecGuard: Benchmarking and Mitigating Over-defense in Prompt Injection Guardrail Models
-Hao Li1,2thanks: These authors contribute equally to this work., Xiaogeng Liu3, 1University of Electronic Science and Technology of China, 2National University of Singapore, 3University of Wisconsin-Madison
+Hao Li1thanks: These authors contribute equally to this work., Xiaogeng Liu2 1Washington University in St. Louis, 2University of Wisconsin-Madison
 1 Introduction
@@
 (2022); Harang (2023a); Perez and Ribeiro (2022); Willison (2022) which can override built-in safeguards and enable the execution of unauthorized actions.
-As the deployment of LLMs continues to expand across industries, the risk posed by prompt injection attacks becomes more pronounced, demanding effective and efficient defense mechanisms.
 Figure 1: Performance comparison of injection detection: We present the average accuracy across benign, malicious, and over-defense cases, plotted against time efficiency.
@@
 These models work by analyzing the semantic meaning of the input data to detect malicious intent before it reaches the LLM.
-Unlike LLM models, prompt guard models are lightweight and computationally efficient, as they do not require the high inference cost associated with LLMs.
+Unlike LLMs, prompt guard models are lightweight and computationally efficient, as they do not require the high inference cost associated with LLMs.
 Additionally, these models operate independently, without the need for victim LLM’s responses, further reducing computation costs compared to approaches like LLM guardrails Inan et al.
@@
 These attributes make prompt guard models adaptable to various environments, and an attractive solution in scenarios where speed and resource optimization are critical.
-We
-Figure 2: Left: PromptGuard Meta (2024) 1, Right: ProtectAIv2 ProtectAI.com (2024) 2, which is current SotA prompt guard model.
-We found that existing prompt guard models have an over-defense issue, i.e., misclassify benign inputs as malicious due to an overreliance on specific trigger words, such as “ignore”.
 Despite these advantages, we find that existing prompt guard models face a critical limitation: the issue of over-defense.
@@
 (2023), where immediate and reliable access is crucial.
+Figure 2: Over-denfese issue in ProtectAIv2 ProtectAI.com (2024) 1, the current SotA prompt guard model.
 To address this issue, we introduce NotInject, an evaluation dataset specifically designed to assess the over-defense issue of existing models.
@@
 Furthermore, our model reaches this performance using a fully open-source dataset, unlike some existing prompt guard models that rely on closed datasets fmops (2024); Meta (2024); ProtectAI.com (2024); LakeraAI (2024a), further promoting a transparency and open-source academic research environment.
-11footnotetext: https://huggingface.co/meta-llama/Prompt-Guard-86M22footnotetext: https://huggingface.co/protectai/deberta-v3-base-prompt-injection-v2
+11footnotetext: https://huggingface.co/protectai/deberta-v3-base-prompt-injection-v2
 2 Related Works
 Prompt injection attacks.
-Prompt injection attacks have become a critical threat to LLMs.
-Since LLMs process inputs in natural language, they often face challenges distinguishing between legitimate user commands and external manipulative inputs, leaving them vulnerable.
-The concept is first identified in research by Perez and Ribeiro (2022), revealing that LLMs could be misled by simple, crafted inputs, resulting in goal hijacking and prompt leakage.
+The concept of prompt injection attacks is first identified in research by Perez and Ribeiro (2022), revealing that LLMs could be misled by simple, crafted inputs, resulting in goal hijacking and prompt leakage.
 Several studies have been proposed Greshake et al.
@@
 Prompt guard models.
-Prompt guard models offer a promising approach to defend against prompt injection attacks by detecting malicious intent in inputs.
+Prompt guard models aim to detect malicious intent in inputs.
 These methods are computationally efficient and, unlike LLM guardrails Inan et al.
 (2023), do not require an additional round of victim LLM inference to generate a response.
-Several prompt guard models have been proposed recently.
-Fmops fmops (2024) is a guardrail trained by FmopsAI, utilizing DistilBERT Sanh et al.
-(2020) as the backbone.
-Deepset Deepset (2024b), PromptGuard Meta (2024), and ProtectAIv2 ProtectAI.com (2024) all employ DeBERTaV3-base He et al.
-(2023) as the backbone.
-Deepset Deepset (2024b) is trained on Deepset-Injection Deepset (2024a), a small open-source dataset with 546 samples.
-PromptGuard is a model trained by Meta with a closed-source dataset.
-ProtectAIv2 is also trained with a closed-source dataset by Protect AI, achieving SotA performance among open-source guard models Deepset (2024b); fmops (2024); Meta (2024); ProtectAI.com (2024) on PINT LakeraAI (2024b).
-Furthermore, LakeraGuard LakeraAI (2024a) is a commercial guard model developed by LakeraAI, with undisclosed model details, such as scale and architecture, and available solely through an API.
-However, most of these prompt guard models fmops (2024); Deepset (2024b); Meta (2024); ProtectAI.com (2024) suffer from over-defense issues, where they rely on shortcuts triggered by certain keywords to make predictions, leading them to incorrectly categorize a benign input that contains the keywords as malicious.
-Moreover, most of these models fmops (2024); Meta (2024); ProtectAI.com (2024); LakeraAI (2024a) are trained using closed-source data with undisclosed implementation or training details.
+Several prompt guard models have been proposed recent including open-sourced Fmops fmops (2024), Deepset Deepset (2024b), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024), and commercial close-sourced LakeraGuard LakeraAI (2024a).
+However, these models are either trained on closed datasets or do not release their training details.
+Most of these prompt guard models fmops (2024); Deepset (2024b); Meta (2024); ProtectAI.com (2024) suffer from over-defense issues, where they rely on shortcuts triggered by certain keywords to make predictions, leading them to incorrectly categorize a benign input that contains the keywords as malicious.
+Moreover, all of these models are trained using closed-source data with undisclosed implementation or training details.
 In this paper, we introduce an over-defense dataset to systematically evaluate the over-defense issue, and propose a novel and completely open-source training approach that achieves SotA performance across diverse benchmarks, including the over-defense dataset.
+Shortcut learning.
+Shortcut learning refers to the common phenomenon across various tasks Niven and Kao (2019); Wang et al.
+(2021); Manjunatha et al.
+(2019), where machine learning models Geirhos et al.
+(2020); Alshemali and Kalita (2020) develop spurious correlations between input features and target labels.
+To address this challenge, a range of mitigation strategies have been proposed, including regularization Wang et al.
+(2021), domain adaptation Liu and Avci (2019), data augmentation Wu et al.
+(2022), multi-task learning Du et al.
+(2021).
+However, recent studies suggest that these techniques may sometimes undermine model performance Khani and Liang (2021), particularly in in-domain distribution (IID) scenarios Mendelson and Belinkov (2021).
+Unlike many previous tasks Niven and Kao (2019); Wang et al.
+(2021); Manjunatha et al.
+(2019); Lin et al.
+(2024), injection attack detection is both a semantic- and pattern-based task, making it more vulnerable to spurious correlations compared to semantic-only tasks.
+Effectively mitigating these correlations without sacrificing performance is a significant challenge for this task.
 Figure 3: Visualization of attention weight.
@@
 3 Over-defense Dataset: NotInject
-Figure 4: The pipeline for constructing NotInject dataset.
+Figure 4: The pipeline for constructing NotInject dataset
 3.1 The Over-defense Issue
@@
 3 with red lines, ProtectAIv2 ProtectAI.com (2024), one of the SotA prompt guard models on PINT benchmark, distributes excessive and unbalanced attention to the word “ignore”, leading it to incorrectly categorize a benign input as malicious.
-To further illustrate, we input a benign sentence containing the word “ignore” into two advanced prompt guard models: PromptGuard Meta (2024) and ProtectAIv2 ProtectAI.com (2024), and present the results in Fig.
-2.
-Both models incorrectly classify the benign instruction containing the word “ignore” as malicious.
+To further illustrate, we input a benign sentence containing the word “ignore” into two advanced prompt guard models, ProtectAIv2 ProtectAI.com (2024) and PromptGuard Meta (2024), with the results separately shown in Fig.
+2 and Fig.
+8.
+Both models incorrectly classifies the benign instruction containing the word “ignore” as malicious.
 In this paper, we define the over-defense issue in prompt guard models as the tendency to predict malicious labels when benign sentences contain certain trigger words commonly used in prompt injection attacks.
@@
 We also visualize the top 20 words identified as trigger words by our method in Fig.
-5.
-Figure 5: The top-20 rank differences between benign and malicious datasets.
-Algorithm 1 Trigger Words Identification
-1:Benign dataset $D_{b}$, Malicious dataset $D_{m}$, Integer $k$
-2:Compute word frequencies in $D_{b}$ to get frequency list $F_{b}$
-3:Compute word frequencies in $D_{m}$ to get frequency list $F_{m}$
-4:Sort $F_{b}$ in descending order to get rank list $R_{b}$
-5:Sort $F_{m}$ in descending order to get rank list $R_{m}$
-6:for all words $w$ in $R_{b}\cup R_{m}$ do
-7:     Compute rank difference $(\Delta r(w)=R_{b}(w)-R_{m}(w))$
-8:     Add $(w,\Delta r(w))$ to list $L$
-9:end for
-10:Sort list $L$ in descending order based on $\Delta f(w)$
-11:return Top $k$ words from list $L$
+9.
 Trigger Words Refinement.
@@
 We first employ LLM to automatically filter the list of candidate words to remove any that may not be pertinent.
-Specifically, We assess the potential harmfulness of words by asking GPT-4o-mini (OpenAI (2024b), we use the 2024-07-18 version by default unless otherwise specified) questions like, “Do you think the word of {word} is especially frequent in malicious or prompt attack scenarios?” 1 Next, we perform a manual verification step, where we review the remaining words after LLM filtering to ensure that any words unrelated to prompt injection attacks are removed.
+Specifically, We assess the potential harmfulness of words by asking GPT-4o-mini (OpenAI (2024b), we use the 2024-07-18 version by default unless otherwise specified) questions like, “Do you think the word of {word} is especially frequent in malicious or prompt attack scenarios?” Next, we perform a manual verification step, where we review the remaining words after LLM filtering to ensure that any words unrelated to prompt injection attacks are removed.
+Specifically, we employ three human evaluators with security expertise to assist in the refinement process.
+Each evaluator independently scores the frequency of each word based on the agreement shown in Fig.
+6.
+The average score for each word is then calculated, and words with an average score above 3 are identified as trigger words.
 Through these two steps, we have compiled a word list with the highest potential for malicious use.
@@
 With a refined list of trigger words, we then generate sample sentences.
-We instruct the GPT-4o-mini to craft new sentences that must include a specified number of the selected trigger words 2.
+We instruct the GPT-4o-mini to craft new sentences that must include a specified number of the selected trigger words.
 Specifically, we select the generated 113 trigger words and create three subsets with distinct difficulty levels, determined by the number of trigger words used in the generation of new sentences.
@@
 To accomplish this, we design a carefully curated prompt (see Appendix.
-C.1) to guide the LLM in generating safe and natural sentences.
+D.1) to guide the LLM in generating safe and natural sentences.
 We then implement a polish process to further ensure the safety of the generated samples.
 In this process, the generated sentences combined with the prompt in Appendix.
-C.2 are re-input into the LLM to identify any potential injection vulnerabilities.
+D.2 are re-input into the LLM to identify any potential injection vulnerabilities.
 Following this, we conduct a manual review to confirm the safety of all sentences and report the error ratio of three subsets in Fig.
-8.
+10.
 This multi-step refinement process guarantees that all generated sentences are harmless.
 The final output forms the proposed NotInject, which contains a total of 339 generated samples with 113 for one-word subset, 113 for two-word subset, and 113 for three-word subset.
-NotInject encompasses a diverse set of topics to enable a thorough evaluation, including common queries from daily life, technique queries (eg., programming, system), virtual creations, and multilingual queries (eg., Chinese, Russian, etc.).
-The detailed category distribution is presented in Table.
-1.
-Category one-word two-word three-word
-Common Queries 58 49 19
-Techniques 16 30 41
-Virtual Creation 14 4 24
-Multilingual 25 30 29
-Table 1: Topic category distribution of NotInject.
-footnotetext: 1, 2 Detailed prompts are in the Appendix. C
+NotInject encompasses a diverse set of topics to enable a thorough evaluation, including common queries from daily life, technique queries (eg., programming, system), virtual creations, and multilingual queries (eg., Chinese, Russian).
+The detailed category distribution is presented in Tab.
+8.
 3.3 Evaluations on NotInject
@@
 Finally, we assess over-defense accuracy, which captures the models’ performance when encountering benign inputs that contain trigger words like “ignore”, using our proposed NotInject dataset.
-Figure 6: Comparison on 3 dimensions of benign, malicious, and over-defense accuracy.
+Figure 5: Comparison of benign, malicious, and over-defense accuracy across various prompt guard solutions.
+InjecGuard significantly outperforms all prior solutions.
+Notably, the open-source models (Deepset, Fmops, PromptGuard, ProtectAIv2) exhibit significant over-defense issues, with over-defense accuracy under 60%, where 50% represents random guessing.
+In addition, although LakeraGuard and LlamaGuard3 demonstrate strong over-defense performance, their effectiveness is still limited by suboptimal malicious accuracy.
 Based on above, we evaluate existing prompt guard models (i.e., Deepset Deepset (2024b), Fmops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024), LakeraGuard LakeraAI (2024a)) and an advanced LLM-based guardrail designed to detect malicious content (i.e., LlamaGuard3 Dubey et al.
@@
 The results are shown in Fig.
-6, where the x, y, and z axes represent malicious accuracy, benign accuracy, and over-defense accuracy, respectively.
+5, where the x, y, and z axes represent malicious accuracy, benign accuracy, and over-defense accuracy, respectively.
 As illustrated, none of the existing prompt guard models (i.e., Deepset Deepset (2024b), Fmops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024)), LakeraGuard LakeraAI (2024a)) nor even the more powerful LLM-based guardrail model (i.e., LlamaGuard3 Dubey et al.
@@
 4 InjecGuard
-In this section, we aim to develop a robust prompt guard model that excels in malicious accuracy, benign accuracy, and over-defense accuracy.
-We also want to emphasize that existing prompt guard models are typically trained in a closed-source manner and disclose their training data.
-Our goal is to train our model using only publicly available data, thereby promoting transparency and facilitating academic research.
-Here, we introduce InjecGuard, a prompt guard model that achieves state-of-the-art performance across various benchmarks.
-To build InjecGuard, we first curate a comprehensive collection of open-source datasets for training and apply data-centric augmentation techniques to address the long-tail problem within the dataset.
-More importantly, we propose a novel training strategy, MOF, which effectively mitigates over-defense issues without relying on specific over-defense datasets.
+We introduce InjecGuard, a prompt guard model with better performance in terms of malicious accuracy, benign accuracy, and over-defense accuracy.
+First, we will introduce data collection and augmentation, followed by our novel training strategy.
 4.1 Data Collection and Augmentation
-Similar to existing models, we first collect a wide range of prompt injection and benign corpus as the training data.
+Almost all existing injection guard solutions (e.g., FMops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024)) have only open-sourced their models, but the datasets and implementation details used for training remain closed-source.
+This means that there is no off-the-shelf public dataset available for training injection guard models.
+Therefore, the first step is to collect a wide range of prompt injection and benign corpus as a basic training dataset.
 Specifically, we select $20$ open-source datasets, including the benign datasets like Alpaca Taori et al.
@@
 Detailed statistics of the training data and prompt for data-centric augmentation are provided in the Appendix.
-A and C.3, respectively.
+A and D.3.
 4.2 Mitigating Over-defense for Free (MOF)
@@
 Ideally, these tokens should all be considered “benign” since they are individual tokens without any intent of prompt injection attack.
-By doing this, we can identify any biased words that are incorrectly predicted as “attack” by the model.
+By doing this, we can identify any biased words that are incorrectly predicted as “attack” by the model trained on basic datasets in Sec.
+4.1.
 We consider the bias toward these tokens as the root cause of the over-defense issue in the model.
 After identifying the biased tokens, we prompt GPT-4o-mini to generate benign data (see Appendix.
-C.1) using random combinations of these tokens (including one-token, two-token, and three-token settings, as mentioned in Sec.
+D.1) using random combinations of these tokens (including one-, two-, and three-token settings, as mentioned in Sec.
 3.2).
-We generate 1,000 benign samples using this method.
+We generate 1,000 benign samples (the scale exploration see Appendix.
+B.2) using this method.
 Afterward, a LLM refinement process similar to the method mentioned in Sec.
-3.2 is employed to ensure the harmlessness of generated data.
+3.2 is employed to ensure the toxicity of generated data.
 Subsequently, we incorporate these generated data into the training data introduced in Sec.
@@
 We also present ablation studies of the aforementioned training design in Tab.
-3.
+2.
 More detailed evaluation results are provided in the following section.
@@
 Evaluation Datasets.
-As mentioned in Sec.
-3.3, we conduct our evaluations based on three aspects: benign accuracy, malicious accuracy, and over-defense accuracy.
-To fully verify the practicality of our proposed InjecGuard, we ensure that all evaluation datasets are not included in the training data.
-Specifically, For evaluating the benign accuracy, we utilize benign data from both the PINT benchmark LakeraAI (2024b) and WildGuard benchmark Han et al.
-(2024).
-For evaluating the malicious accuracy, We use attack data from both the PINT benchmark LakeraAI (2024b) and BIPIA datasets Yi et al.
-(2023).
-For evaluating the over-defense accuracy, we use the proposed NotInject dataset.
-Metrics.
-Since prompt guard models function as text classification systems that predict whether an input text is benign or malicious, we evaluate their performance using Accuracy, calculated as the proportion of correct predictions over the total number of test cases in the evaluation dataset:
-Accuracy=Number of Correct PredictionsTotal Number of Test $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Test Cases}}$ (1)
-Additionally, we report the computational overhead in terms of Giga Floating Point Operations (GFLOPs), which quantifies the total number of floating-point operations required during inference.
-GFLOPs provide an estimate of the computational resources needed by the model.
-We also measure the inference time to assess the computational overhead introduced by the models.
+We evaluate (1) benign accuracy, which utilizes benign data from both the PINT benchmark LakeraAI (2024b) and WildGuard benchmark Han et al.
+(2024); (2) malicious accuracy, which uses attack data from both the PINT benchmark LakeraAI (2024b) and BIPIA datasets Yi et al.
+(2023) and over-defense on our proposed NotInject dataset.
+[…]
 Training Details of InjecGuard.
@@
 5.2 Main Results
-We compare our method with other baselines in terms of performance and overhead. The results are shown in Tab. 2 (detailed results for each subsets deferred to Appendix. B.1).
+We compare our method with other baselines in terms of performance and overhead. The results are shown in Tab. 1 (detailed results for each subset deferred to Appendix. B.1).
 Performance Comparison.
 Our InjecGuard demonstrates superior performance compared to existing prompt guard models and even rivals commercial LLMs like GPT-4o.
-Specifically, InjecGuard achieves an average accuracy of 83.48%, the best-performing commercial prompt guard model, by 6.25%, and exceeding the top open-source model, ProtectAIv2, by 30.83%.
+Specifically, InjecGuard achieves an average accuracy of 83.48%, the best-performing commercial prompt guard model, LakeraGuard, by 6.25%, and exceeding the top open-source model, ProtectAIv2, by 30.83%.
 Our model excels across all evaluation categories, attaining an over-defense accuracy of 87.32%, benign accuracy of 85.74%, and malicious accuracy of 77.39%.
@@
 Notably, according to Tab.
-2, the Efficiency (Performance/Inference time) of our model is 503 times higher than that of GPT-4o.
+1, the Efficiency (Performance/Inference time) of our model is 503 times higher than that of GPT-4o.
 The results also prove that existing prompt guard models suffer from overwhelming issues with over-defense, as reflected by their low over-defense accuracy scores.
@@
 Our model, InjecGuard, excels in both computational efficiency and performance compared to existing baseline models.
-As shown in Table 2, InjecGuard  achieves an average accuracy of 83.48% while maintaining a GFLOPS count of 60.45 and an inference time of only 15.34 milliseconds.
+As shown in Tab.
+1, InjecGuard  achieves an average accuracy of 83.48% while maintaining a GFLOPS count of 60.45 and an inference time of only 15.34 milliseconds.
 With an efficiency score of 5.44 (performance divided by inference time), InjecGuard delivers robust results swiftly and resource-effectively, making it ideal for environments prioritizing both accuracy and efficiency.
@@
 Similarly, models such as Llama-2-chat and LlamaGuard3 require thousands of GFLOPS and longer inference times, resulting in much lower efficiency.
-This disparity highlights the impracticality of deploying these LLMs in scenarios where computational resources and quick response times are essential.
-Therefore, InjecGuard presents a highly effective and efficient alternative for real-world applications demanding both high accuracy and operational efficiency.
-Training Design Over-defense (%) Benign (%) Malicious (%) Average (%)
-Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
-   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
-Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
-   + MOF with Retraining from Scratch (Sec. 4.2) 89.38 84.73 71.57 81.89
-Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
-   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
-   + MOF with Finetuning (Sec. 4.2) 68.14 82.11 74.81 75.02
-Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
-   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
-   + MOF with Retraining from Scratch (Sec. 4.2) 87.32 85.74 77.39 83.48
-Table 3: Ablation study of each training design introduced in Sec. 4.
+Training Design Overdef. Benign Mal. Avg.
+Basic dataset training 75.22 78.53 70.17 74.64
++ Data-centric augment 64.31 81.36 75.95 73.87
+Basic dataset training 75.22 78.53 70.17 74.64
++ MOF scratch retrain 89.38 84.73 71.57 81.89
+Basic dataset training 75.22 78.53 70.17 74.64
++ Data-centric augment 64.31 81.36 75.95 73.87
++ MOF with finetuning 68.14 82.11 74.81 75.02
+Basic dataset training 75.22 78.53 70.17 74.64
++ Data-centric augment 64.31 81.36 75.95 73.87
++ MOF scratch retrain 87.32 85.74 77.39 83.48
+Table 2: Ablation study of each training design.
 5.3 Ablation Studies
@@
 2 offers key insights into the effects of our training components on InjecGuard performance.
-Starting with the Basic Dataset, the model establishes a baseline average accuracy of 74.64%, with over-defense, benign, and malicious accuracies of 75.22%, 78.53%, and 70.17%, respectively.
+Starting with the Basic Dataset (Sec.
+4.1), the model establishes a baseline average accuracy of 74.64%, with over-defense, benign, and malicious accuracies of 75.22%, 78.53%, and 70.17%, respectively.
 Introducing Data-centric Augmentation alone leads to a noticeable improvement in benign accuracy (from 78.53% to 81.36%) and malicious accuracy (from 70.17% to 75.95%), enhancing the model’s ability to correctly classify both benign and malicious inputs.
@@
 The ablation study demonstrates that data-centric augmentation alone can improve certain aspects of model performance, and the integration of MOF techniques is essential to counterbalance the increased over-defense, thereby ensuring that InjecGuard achieves optimal and balanced performance across all metrics.
+Method Over-defense (%) Malicious (%)
+InjecGuard (w/o MOF) 64.31 75.95
++ shortcut mitigation 86.73 65.53
++ MOF 87.32 77.39
+Table 3: Comparison with shortcut mitigation method.
+5.4 Comparison with Shortcut Mitigation
+In Sec.
+2, we discuss the susceptibility of injection attack detection tasks to spurious correlations.
+These correlations emerge from the interplay between semantic understanding and pattern recognition requirements, potentially challenging conventional shortcut mitigation methods.
+To further investigate the effectiveness of MOF compared to typical shortcut mitigation methods, we implement a representative shortcut mitigation approach Wang et al.
+(2021) to InjecGuard without MOF.
+The results presented in Tab.
+3 show that while the shortcut mitigation method Wang et al.
+(2021) improves over-defense performance, it leads to a significant 10.42% decline in malicious performance, a well-known issue in current shortcut mitigation techniques Khani and Liang (2021); Mendelson and Belinkov (2021).
+In contrast, MOF not only achieves superior overall defense performance (87.32%) but also improves malicious performance to 77.39%, demonstrating the advantages of MOF over existing shortcut mitigation methods in this pattern-based task.
 6 Conclusions
-In this study, we delved into the over-defense phenomenon observed in existing injection-based prompt guard models.
-To address this issue, we proposed a novel over-defense data sampling framework and introduced the NotInject benchmark to systematically evaluate the extent of over-defense in target prompt guard models.
-Furthermore, we presented InjecAgent, an advanced prompt guard model trained using a novel robust training strategy, MOF, designed to effectively mitigate model bias toward specific words.
-Our experimental results on various benchmark demonstrate that InjecAgent significantly outperforms existing prompt guard models across multiple dimensions, with substantial improvements in performance and robustness.
+In this paper, we demonstrated the over-defense phenomenon in existing prompt guard models.
+To address this, we introduced the NotInject benchmark to evaluate the extent of over-defense.
+Furthermore, we presented InjecGuard, a prompt guard model that can significantly outperform existing models in both performance and robustness.
 To the best of our knowledge, this is the first work to provide a fully open-source prompt guard model against injection, including the training dataset, strategies, code, and model.
@@
 Table 7: Full results of comparison between existing injection guardrails.
+Category one-word two-word three-word
+Common Queries 58 49 19
+Techniques 16 30 41
+Virtual Creation 14 4 24
+Multilingual 25 30 29
+Table 8: Topic category distribution of NotInject.
 Appendix B Additional Experimental Results
@@
 In Tab.
-2, we have illustrated the comprehensive results on different dimensions, such as over-defense, benign, and malicious.
+1, we have illustrated the comprehensive results on different dimensions, such as over-defense, benign, and malicious.
 In this section, we present detailed results for each evaluation benchmark across all dimensions, including our NotInject, Wildguard Han et al.
@@
 7.
-B.2 Visualization of NotInject dataset
+Number over-defense benign malicious average
+500 74.63 82.19 80.55 79.12
+1,000 87.32 85.74 77.39 83.48
+2,000 92.63 85.99 71.12 83.24
+Table 9: Ablation study of MOF sampling scale.
+B.2 Ablation study of MOF sampling scale
+In Sec.
+4.2, MOF generates 1,000 benign samples containing biased tokens, which are incorporated into the training process to reduce the model’s over-defense degree.
+To explore the optimal scale for generating benign samples, we conduct experiments using MOF to generate samples at varying scales and evaluate the performance of InjecGuard trained on these datasets.
+The results shown in Tab.
+9 reveal a clear trade-off between over-defense and malicious accuracy as the number of generated samples increases.
+Notably, training with 1,000 generated samples achieves the best balance, resulting in the highest average accuracy of 83.48%.
+We thereby select 1,000 as the sampling scale.
+B.3 The effect of MOF on general over-defense scenarios
+Although MOF primarily focuses on non-trigger word scenarios, we further explore the effectiveness of MOF in general over-defense scenarios.
+For the experimental details, since there is no existing benchmark for non-trigger-word-based over-defense, we adopt the hard-negative dataset from the PINT LakeraAI (2024b) benchmark as a substitute, which can somewhat reflect the model’s performance in non-trigger-word-based over-defense scenarios.
+The results in Tab.
+10 clearly show that incorporating MOF improves performance on the hard-negative dataset, demonstrating its effectiveness in mitigating over-defense in non-trigger-word-based scenarios.
+The accuracy of 91.15% further highlights the robustness of InjecGuard across general over-defense scenarios.
+Method hard negative accuracy
+InjecGuard (w/o MOF) 87.86
+InjecGuard (w/ MOF) 91.15
+Table 10: The effect of MOF on general over-defense scenarios.
+Appendix C Qualitative Analysis
+C.1 Visualization of NotInject dataset
 To further investigate the advantages of our method on confronting over-defense, we select the benign sentence from our over-defense dataset and perform a visualization to conduct qualitative analysis for model predictions.
@@
 In contrast, our InjecGuard accurately classifies the input as safe, highlighting the efficacy of the over-defense dataset in evaluating over-defense tendencies and demonstrating the robustness of the proposed InjecGuard.
+C.2 The robustness of MOF when incorrectly identified words
+In this section, we explore the robustness of the MOF when trigger words are incorrectly identified and assess whether such misidentifications could harm model performance.
+In fact, incorrectly identified words do not have a detrimental effect on the model.
+This is because these misidentified words are used to construct benign sentences, which do not introduce noisy labels or degrade the model’s learning process.
+For example, if an unbiased word like “book” is mistakenly identified as biased, and we construct a sentence such as “Can you recommend some history books for me?” adding this sentence to the training dataset would merely serve as a benign augmentation.
+While this might slightly reduce the degree of improvement in over-defense performance, it would not negatively impact the model’s basic performance.
+Figure 6: The human curation agreement in trigger word refinement process.
 Figure 7: Different guard models’ prediction visualization.
-Appendix C Prompts
+Figure 8: Over-denfese issue in PromptGuard Meta (2024) 1.
+Figure 9: The top-20 rank differences between benign and malicious datasets.
+Figure 10: The error ratio of three subsets after LLM generation.
+Algorithm 1 Trigger Words Identification
+1:Benign dataset $D_{b}$, Malicious dataset $D_{m}$, Integer $k$
+2:Compute word frequencies in $D_{b}$ to get frequency list $F_{b}$
+3:Compute word frequencies in $D_{m}$ to get frequency list $F_{m}$
+4:Sort $F_{b}$ in descending order to get rank list $R_{b}$
+5:Sort $F_{m}$ in descending order to get rank list $R_{m}$
+6:for all words $w$ in $R_{b}\cup R_{m}$ do
+7:     Compute rank difference $(\Delta r(w)=R_{b}(w)-R_{m}(w))$
+8:     Add $(w,\Delta r(w))$ to list $L$
+9:end for
+10:Sort list $L$ in descending order based on $\Delta f(w)$
+11:return Top $k$ words from list $L$
+Appendix D Prompts
 In this section, we illustrate the prompts used in our method.
-C.1 Word-based Generation Prompt
+D.1 Word-based Generation Prompt
 In Sec.
@@
 The prompts used in this process are detailed in Fig.
-9.
-C.2 Refinement Prompt
+11.
+D.2 Refinement Prompt
 In Sec.
@@
 The detailed prompts are provided in Fig.
-10.
-C.3 Long-tail Format Augmentation Prompt
+12.
+D.3 Long-tail Format Augmentation Prompt
 In Sec.
@@
 The detailed augmentation prompt is illustrated in Fig.
-11.
-Figure 8: The error ratio of three subsets after LLM generation.
+13.
 Figure 11: Word-based Generation Prompt.
@@
 Figure 13: LLM Augmentation Prompt.
+11footnotetext: https://huggingface.co/meta-llama/Prompt-Guard-86M